# Backstepping for Partial Differential Equations

### v2

@@ Abstract
Systems modeled by partial differential equations (PDEs) are at least as ubiquitous as those by nature finite-dimensional and modeled by ordinary differential equations (ODEs). And yet, systematic and readily usable methodologies, for such a significant portion of real systems, have been historically scarce. Around the year 2000, the backstepping approach to PDE control began to offer not only a less abstract alternative to PDE control techniques replicating optimal and spectrum assignment techniques of the 1960s, but also enabled the methodologies of adaptive and nonlinear control, matured in the 1980s and 1990s, to be extended from ODEs to PDEs, allowing feedback synthesis for systems that are uncertain, nonlinear, and infinite-dimensional. The PDE backstepping literature has since grown to hundreds of papers and nearly a dozen books. This survey aims to facilitate the entry into this thriving area of overwhelming size and topical diversity. Designs of controllers and observers, for parabolic, hyperbolic, and other classes of PDEs, in one and more dimensions, with nonlinear, adaptive, sampled-data, and event-triggered extensions, are covered in the survey. The lifeblood of control are technology and physics. The survey places a particular emphasis on applications that have motivated the development of the theory and which have benefited from the theory and designs: flows, flexible structures, materials, thermal and chemically reacting dynamics, energy (from oil drilling to batteries and magnetic confinement fusions), and vehicles.
@@ 1 Introduction
More specifically, the control of systems governed by PDEs has applications in many branches of science and engineering. The ability to effectively control these systems allows optimizing performance, ensuring safety, and achieving desired behaviors in a wide range of applications.
In this context, backstepping for PDEs has emerged as a powerful and versatile methodology for crafting boundary controllers and observers in the infinite-dimensional setting. Over the past two decades, this constructive approach has been successfully applied to a wide spectrum of PDE systems, as evidenced by the burgeoning number of publications in this domain, comprising not only several hundreds of papers but also almost a dozen of books, see e.g. [270, 170, 48, 179, 215, 253, 284, 302, 13], some with over a thousand citations. The only (then-)general textbook was published in 2008 [194] and in the last 15 years the field has at least quintupled in size and breadth. The proliferation of research has reached a point where it is increasingly challenging for novices to navigate and digest the extensive literature.
@@ 1.1 Origins of PDE backstepping
The roots of the method can be traced back to the 1970s and 1980s, when Volterra-type integral transforms were employed to investigate solutions and controllability properties of PDEs in a handful of lesser-known works [81, 242]. However, it was not until the early 2000s that these transformations were independently rediscovered and leveraged for feedback design and stabilization.
Backstepping, as a control design methodology, first gained prominence in the context of finite-dimensional systems. The origins of the backstepping idea are difficult to trace, because of its simultaneous appearance, often implicit, in the works of Tsinias [259], Koditschek [167], Byrnes and Isidori [70], and Sontag and Sussmann [254]. Backstepping was initially conceived for nonlinear systems with an integrator at the input, such as in robotic systems where the kinematics are separated from the input torque by angular velocity, which integrates in time the torque and other sources of angular acceleration. The recursive application of backstepping led to its generalization to strict-feedback nonlinear systems, which are of the form . The term (strict)-feedback refers to the “recirculation of the states of the integrator chain into the -th equation through the nonlinearity , which is free of the dependencies on the ‘feedforward’ state variables through . The book [187], often regarded as the definitive backstepping reference, has garnered thousands of citations. The popularity of backstepping in the finite-dimensional setting can be attributed to several factors: the ubiquity of strict-feedback systems in applications, the constructive nature of the method, and its ability to yield explicit control laws—features seldom found in nonlinear control. Moreover, the approach can be extended to design adaptive controllers [187].
@@ 1.2 Alternative methods for PDE Control
Another traditional approach is the spectral decomposition method, which exploits the eigenfunction expansion of the PDE operator to transform the system into an infinite series of ODEs [196, 40]. This allows for the application of finite-dimensional control techniques to each mode separately. For instance, pole-placement theorems with spectral decomposition of Sturm-Liouville operators can be used to design finite-dimensional output feedback controllers [202]. However, this approach is limited to systems with simple geometries and boundary conditions, and it may suffer from slow convergence and numerical instabilities. In addition it is not clear how such methods could be made adaptive, when the PDE’s parameters are unknown, since the dimension of the reduced model would vary in real time, as the parameter estimate varies.
@@
Model reference adaptive control (MRAC) approaches have also been extended to distributed parameter systems. For example, Böhm et al. [57] developed an infinite-dimensional MRAC framework for a broad class of nonlinear distributed parameter systems. While the approach provides theoretical results on well-posedness, stability, and parameter convergence along with finite-dimensional approximation schemes, it requires full state measurements and fully distributed (in-domain) control inputs, which are seldom found in practical applications.
@@ 2 A primer in backstepping for PDEs
This section explores the backstepping methodology for PDEs in more detail using the aforementioned reaction-diffusion designs as a basic paradigmatic example. After introducing some notations and basic PDE mathematics, we explore the selection of target systems, transformations, and kernel equations, culminating in the practical application of these concepts to ensure system stability. The discussion also covers with the design of observers and output feedback control laws. We finish the section with other basic early backstepping designs. This gradual exposition lays the foundation for understanding how backstepping principles apply to more complex systems, explored in later sections.
2.1 Mathematical Preliminaries: Functional Spaces, Norms, Inequalities and Stability
@@
Of the spaces, is the most used in backstepping, as it is distinguished by being a Hilbert space (a complete inner product vector space), equipped with the inner product , facilitating the use of geometric and projection methods. However, properties of functions, such as continuity, boundedness, or differentiability, are not guaranteed within spaces alone. For this purpose, functional spaces that account for spatial derivatives are necessary; these are known as Sobolev spaces. We define two primary Sobolev spaces, and , with norms defined as and . The spaces and are composed of all functions, with the corresponding norm being finite. They are both Hilbert spaces equipped with inner products. In this survey, we mostly use the , , and spaces and norms.
@@
The constants and in (1) are respectively referred to as the overshoot coefficient and the decay rate. Similar statements may be posed for other norms, and combined estimates are usually necessary for systems with more than one state. An example on how to derive estimates like (1) is presented next in Section 2.2.
@@ 2.2 Reaction-diffusion PDEs and their stability
where represents the state, for and , and denotes a possible actuation variable. In . (2), the diffusion coefficient is assumed to be positive, and the reaction term is considered continuous. We adopt a Dirichlet-type left boundary condition, , although Neumann () or Robin ( for some ) boundary conditions can also be utilized. The Dirichlet actuation in Eq. (2) could also alternatively be replaced with Neumann actuation. In this context, the most appropriate spaces333See [135] for details of how a linear 1-D parabolic PDE with Dirichlet boundary conditions is well-posed in and . In this survey, the well-posedness of the PDEs under study is assumed, given that it represents a complex subject on its own. to operate in are or . Considering the former, assume initial conditions with .
Consider first the open loop system by setting in (2). Let represent the upper bound of within the interval . When , the equilibrium state of (2) can be shown exponentially stable in open loop, and an estimate similar to (1) can be derived, as demonstrated next. Consider the following Lyapunov functional, equivalent to the squared norm:
@@ 2.3 Backstepping transformations and target systems
Continuing our example, we turn our attention to posing a transformation that maps system (2) onto the selected target system (7). To fully grasp the appropriate structure for such a transformation, it is instructive to revisit the origins of backstepping briefly. Initially developed for finite-dimensional strict-feedback systems as explained in Section 1.1, backstepping addressed configurations where a sequential chain of integrators ends with the control variable in the final equation. These systems display a “lower-triangular” structure, wherein the right-hand side terms of successive equations progressively depend on the current and all preceding states, thus establishing a cascading pattern. The successful application of backstepping to lower-triangular ODEs inspired its adaptation to PDEs, notably reaction-diffusion systems as exemplified by (2), which manifest “diagonal” dynamics. In these dynamics, specific terms, such as the reaction term or the time derivative of the state , are influenced solely by the local state , although the diffusion term extends this dependence to the infinitesimal neighborhood. Moreover, additional terms might emulate a strict-feedback structure, for instance, and , thereby introducing a notion of “spatial causality” where the temporal evolution at any point is contingent on both the current state and those preceding it, i.e., for . It should be noted, however, that this structure does not implies a finite speed of propagation; in a reaction-diffusion system, disturbances anywhere instantaneously influence the entire domain. This is analogous to a chain of integrators of any length, where perturbations instantaneously impact all states.
@@
This transformation is henceforth designated as the backstepping transformation, with its integral kernel termed the backstepping kernel. It is characterized by a Volterra-type integral, and is therefore invertible under relatively lenient conditions on the kernel , a subject elaborated upon in Section 2.6.
@@ 2.4 Kernel equations: derivation and well-posedness
Once the backstepping transformation and target system have been chosen, the next step is to find the conditions that the backstepping kernel needs to verify, ensuring that the original system is mapped into the desired target system. These conditions take the form of (usually) hyperbolic PDEs, known as the kernel equations, and their well-posedness and solvability (or lack thereof) is what ultimately determines if the target system and transformation were correctly chosen. The way to derive the kernel equations is by substituting the transformation (8) into the target system (7) and eliminating . Then, integrating by parts, we get a weak formulation of a PDE, which can then be written in standard form. In the procedure, the (uncontrolled) boundary condition is also taken into account.
@@
defines a unique continuous solution for the integral form of the kernel equations and therefore there exists a unique continuous solving (10)–(12) in . This procedure also can be used to show that given that is continuous.
@@ 2.6 Invertibility of the transformation and closed loop stability
Equation (20) does not depend on the original or target systems and allows to compute from . An alternative approach is to derive a PDE equation for the kernel, which is done in a similar fashion to the kernel equation. Note, however, that (at least in this design) the inverse transformation is not used in the implementation of backstepping; only its existence and properties are required for the proof of closed-loop stability. In particular, the norm of this inverse transformation appears in the stability estimate as seen below in (24).
@@ 2.9 Extensions to other one-dimensional single-variable systems
After the initial results for parabolic equations was obtained, it was quickly recognized that the methodology presented in the previous sections could be extended to various other PDE systems. Here, several important extensions focusing on 1D systems described by a single PDE are explored. We first discuss designs for a first-order hyperbolic PDE, which models, for example, traffic flows. We then present the extensions to the wave equation and flexible structure models (shear beam and Euler-Bernoulli beam). We mention how backstepping can be applied to complex-valued PDEs such as the Schrodinger and Ginzburg-Landau equations. We finish with an example where backstepping is used to prove null controllability for a Korteweg-de Vries equation.
@@ 2.9.1 First-order hyperbolic PDEs
In [193], Krstic and Smyshlyaev extended the backstepping method to solve the stabilization problem for a class of first-order hyperbolic PDEs. The target systems in [193] are chosen to be pure delays, ensuring convergence to zero in finite time.
@@ 2.9.2 Wave equation and flexible structures
The extension of backstepping to wave equations and flexible structure models proved initially challenging due to the lack of inherent damping in their dynamics. While transformations to cancel destabilizing terms in the equation could be found, it was not clear how to ensure the stability of the resulting target systems. The crucial step in extending backstepping to these systems involved identifying doable target systems that incorporated damping terms.
Thus, a problem for a wave equation with anti-damping on the uncontrolled boundary was addressed in [252]. In this scenario, the anti-damping term acted through the boundary condition, making the system unstable with infinitely many eigenvalues in the right-half plane. Another example considers an output-feedback problem with instability at its free end and control on the opposite end [185]. A similar problem of boundary stabilization for a one-dimensional wave equation with an internal spatially varying antidamping term was tackled in [244]. This antidamping term made the uncontrolled system unstable, with all of its eigenvalues located in the right-half plane. The proposed design used a novel backstepping transformation with a 2x2 structure, mapping the unstable system into a stable target system with a desired decay rate. The control gains were expressed in closed form for systems with constant parameters. Other backstepping designs on the wave equation include [310, 291]. Although these backstepping controllers may have limited practical applicability due to their poor delay-robustness margins-resulting from the infinite number of unstable poles in open-loop -the underlying theoretical developments have been essential for the development of the backstepping methodology.
For beams, a pioneering work in this area was [190], where they designed boundary controllers and observers for the wave equation and the Timoshenko beam model with Kelvin-Voigt damping. In contrast to the reaction-diffusion equation, where one could simply cancel the potentially destabilizing reaction term, the target systems in [190] contained damping terms arising from Kelvin-Voigt damping (a type of material damping that introduces a strain-rate term in the constitutive equation). This contribution laid the foundation for subsequent studies on the shear beam [184] and the Euler–Bernoulli beam [252].

### v3

@@ Abstract
Systems modeled by partial differential equations (PDEs) are at least as ubiquitous as those by nature finite-dimensional and modeled by ordinary differential equations (ODEs). And yet, systematic and readily usable methodologies, for such a significant portion of real systems, have been historically scarce. Around the year 2000, the backstepping approach to PDE control began to offer not only a less abstract alternative to PDE control techniques replicating optimal and spectrum assignment techniques of the 1960s, but also enabled the methodologies of adaptive and nonlinear control, matured in the 1980s and 1990s, to be extended from ODEs to PDEs, allowing feedback synthesis for systems that are uncertain, nonlinear, and infinite-dimensional. The PDE backstepping literature has since grown to hundreds of papers and nearly a dozen books. This survey aims to facilitate the entry into this thriving area of overwhelming size and topical diversity. Designs of controllers and observers, for parabolic, hyperbolic, and other classes of PDEs, in one or more dimensions, with nonlinear, adaptive, sampled-data, and event-triggered extensions, are covered in the survey. The lifeblood of control are technology and physics. The survey places a particular emphasis on applications that have motivated the development of the theory and which have benefited from the theory and designs: flows, flexible structures, materials, thermal and chemically reacting dynamics, energy (from oil drilling to batteries and magnetic confinement fusion), and vehicles.
@@ 1 Introduction
More specifically, the control of systems governed by PDEs has applications in many branches of science and engineering. The ability to effectively control these systems allows for optimizing performance, ensuring safety, and achieving desired behaviors in a wide range of applications.
In this context, backstepping for PDEs has emerged as a powerful and versatile methodology for crafting boundary controllers and observers in the infinite-dimensional setting. Over the past two decades, this constructive approach has been successfully applied to a wide spectrum of PDE systems, as evidenced by the burgeoning number of publications in this domain, comprising not only several hundreds of papers but also almost a dozen of books, see e.g. [273, 171, 48, 180, 216, 255, 288, 306, 13], some with over a thousand citations. The only (then-)general textbook was published in 2008 [195], and in the last 15 years, the field has at least quintupled in size and breadth. The proliferation of research has reached a point where it is increasingly challenging for novices to navigate and digest the extensive literature.
@@ 1.1 Origins of PDE backstepping
The roots of the method can be traced back to the 1970s and 1980s when Volterra-type integral transforms were employed to investigate solutions and controllability properties of PDEs in a handful of lesser-known works [82, 244]. However, it was not until the early 2000s that these transformations were independently rediscovered and leveraged for feedback design and stabilization.
Backstepping, as a control design methodology, first gained prominence in the context of finite-dimensional systems. The origins of the backstepping idea are difficult to trace, because of its simultaneous appearance, often implicit, in the works of Tsinias [261], Koditschek [168], Byrnes and Isidori [71], and Sontag and Sussmann [256]. Backstepping was initially conceived for nonlinear systems with an integrator at the input, such as in robotic systems where the kinematics are separated from the input torque by angular velocity, which integrates in time the torque and other sources of angular acceleration. The recursive application of backstepping led to its generalization to strict-feedback nonlinear systems, which are of the form . The term (strict)-feedback refers to the “recirculation” of the states of the integrator chain into the -th equation through the nonlinearity , which is free of the dependencies on the ‘feedforward’ state variables through . The book [188], often regarded as the definitive backstepping reference, has garnered thousands of citations. The popularity of backstepping in the finite-dimensional setting can be attributed to several factors: the ubiquity of strict feedback systems in applications, the constructive nature of the method, and its ability to yield explicit control laws—features seldom found in nonlinear control. Moreover, the approach can be extended to design adaptive controllers [188].
@@ 1.2 Alternative methods for PDE Control
Another traditional approach is the spectral decomposition method, which exploits the eigenfunction expansion of the PDE operator to transform the system into an infinite series of ODEs [197, 40]. This allows for the application of finite-dimensional control techniques to each mode separately. For instance, pole-placement theorems with spectral decomposition of Sturm-Liouville operators can be used to design finite-dimensional output feedback controllers [203]. However, this approach is limited to systems with simple geometries and boundary conditions, and it may suffer from slow convergence and numerical instabilities. In addition, it is not clear how such methods could be made adaptive when the PDE’s parameters are unknown since the dimension of the reduced model would vary in real-time, as the parameter estimate varies.
@@
Model reference adaptive control (MRAC) approaches have also been extended to distributed parameter systems. For example, Böhm et al. [57] developed an infinite-dimensional MRAC framework for a broad class of nonlinear distributed parameter systems. While the approach provides theoretical results on well-posedness, stability, and parameter convergence along with finite-dimensional approximation schemes, it requires full-state measurements and fully distributed (in-domain) control inputs, which are seldom found in practical applications.
@@ 2 A primer in backstepping for PDEs
This section explores the backstepping methodology for PDEs in more detail using the aforementioned reaction-diffusion designs as a basic paradigmatic example. After introducing some notations and basic PDE mathematics, we explore the selection of target systems, transformations, and kernel equations, culminating in the practical application of these concepts to ensure system stability. The discussion also covers the design of observers and output feedback control laws. We finish the section with other basic early backstepping designs. This gradual exposition lays the foundation for understanding how backstepping principles apply to more complex systems, explored in later sections.
2.1 Mathematical Preliminaries: Functional Spaces, Norms, Inequalities, and Stability
@@
Of the spaces, is the most used in backstepping, as it is distinguished by being a Hilbert space (a complete inner product vector space), equipped with the inner product , facilitating the use of geometric and projection methods. However, properties of functions, such as continuity, boundedness, or differentiability, are not guaranteed within spaces alone. For this purpose, functional spaces that account for spatial derivatives are necessary; these are known as Sobolev spaces. We define two primary Sobolev spaces, and , with norms defined as and . The spaces and are composed of all functions, with the corresponding norm being finite. They are both Hilbert spaces equipped with inner products. In this survey, we mostly use the , , , and spaces and norms.
@@
The constants and in (1) are respectively referred to as the overshoot coefficient and the decay rate. Similar statements may be posed for other norms, and combined estimates are usually necessary for systems with more than one state. An example of how to derive estimates like (1) is presented next in Section 2.2.
@@ 2.2 Reaction-diffusion PDEs and their stability
where represents the state, for and , and denotes a possible actuation variable. In (2), the diffusion coefficient is assumed to be positive, and the reaction term is considered continuous. We adopt a Dirichlet-type left boundary condition, , although Neumann () or Robin ( for some ) boundary conditions can also be utilized. The Dirichlet actuation in Eq. (2) could also alternatively be replaced with Neumann actuation. In this context, the most appropriate spaces333See [136] for details of how a linear 1-D parabolic PDE with Dirichlet boundary conditions is well-posed in and . In this survey, the well-posedness of the PDEs under study is assumed, given that it represents a complex subject on its own. to operate in are or . Considering the former, assume initial conditions with .
Consider first the open loop system by setting in (2). Let represent the upper bound of within the interval . When , the equilibrium state of (2) can be shown to be exponentially stable in open loop, and an estimate similar to (1) can be derived, as demonstrated next. Consider the following Lyapunov functional, equivalent to the squared norm:
@@ 2.3 Backstepping transformations and target systems
Continuing our example, we turn our attention to posing a transformation that maps system (2) onto the selected target system (7). To fully grasp the appropriate structure for such a transformation, it is instructive to revisit the origins of backstepping briefly. Initially developed for finite-dimensional strict-feedback systems as explained in Section 1.1, backstepping addressed configurations where a sequential chain of integrators ends with the control variable in the final equation. These systems display a “lower-triangular” structure, wherein the right-hand side terms of successive equations progressively depend on the current and all preceding states, thus establishing a cascading pattern. The successful application of backstepping to lower-triangular ODEs inspired its adaptation to PDEs, notably reaction-diffusion systems as exemplified by (2), which manifest “diagonal” dynamics. In these dynamics, specific terms, such as the reaction term or the time derivative of the state , are influenced solely by the local state , although the diffusion term extends this dependence to the infinitesimal neighborhood. Moreover, additional terms might emulate a strict-feedback structure, for instance, and , thereby introducing a notion of “spatial causality” where the temporal evolution at any point is contingent on both the current state and those preceding it, i.e., for . It should be noted, however, that this structure does not imply a finite speed of propagation; in a reaction-diffusion system, disturbances anywhere instantaneously influence the entire domain. This is analogous to a chain of integrators of any length, where perturbations instantaneously impact all states.
@@
This transformation is henceforth designated as the backstepping transformation, with its integral kernel termed the backstepping kernel. It is characterized by a Volterra-type integral and is therefore invertible under relatively lenient conditions on the kernel , a subject elaborated upon in Section 2.6.
@@ 2.4 Kernel equations: derivation and well-posedness
Once the backstepping transformation and target system have been chosen, the next step is to find the conditions that the backstepping kernel needs to verify, ensuring that the original system is mapped into the desired target system. These conditions take the form of (usually) hyperbolic PDEs, known as the kernel equations, and their well-posedness and solvability (or lack thereof) are what ultimately determine if the target system and transformation were correctly chosen. The way to derive the kernel equations is by substituting the transformation (8) into the target system (7) and eliminating . Then, integrating by parts, we get a weak formulation of a PDE, which can then be written in standard form. In the procedure, the (uncontrolled) boundary condition is also taken into account.
@@
defines a unique continuous solution for the integral form of the kernel equations, and therefore, there exists a unique continuous solving (10)–(12) in . This procedure also can be used to show that given that is continuous.
@@ 2.6 Invertibility of the transformation and closed loop stability
Equation (20) does not depend on the original or target systems and allows to compute from . An alternative approach is to derive a PDE equation for the kernel, which is done in a similar fashion to the kernel equation. Care should be taken in more general designs (e.g., those including extra degrees of freedom in the kernel equations) to make sure the kernel thus obtained is indeed the inverse of the kernel. Note, however, that (at least in this design) the inverse transformation is not used in the implementation of backstepping; only its existence and properties are required for the proof of closed-loop stability. In particular, the norm of this inverse transformation appears in the stability estimate as seen below in (24).
@@ 2.9 Extensions to other one-dimensional single-variable systems
After the initial results for parabolic equations were obtained, it was quickly recognized that the methodology presented in the previous sections could be extended to various other PDE systems. Here, several important extensions focusing on 1D systems described by a single PDE are explored. We first discuss designs for a first-order hyperbolic PDE, which models, for example, traffic flows. We then present the extensions to the wave equation and flexible structure models (shear beam and Euler-Bernoulli beam). We mention how backstepping can be applied to complex-valued PDEs such as the Schrödinger and Ginzburg-Landau equations. We finish with an example where backstepping is used to prove null controllability for a Korteweg-de Vries equation.
@@ 2.9.1 First-order hyperbolic PDEs
In [194], Krstic and Smyshlyaev extended the backstepping method to solve the stabilization problem for a class of first-order hyperbolic PDEs. The target systems in this design are chosen to be pure delays, ensuring convergence to zero in finite time.
@@ 2.9.2 Wave equation and flexible structures
The extension of backstepping to wave equations and flexible structure models proved initially challenging due to the lack of inherent damping in their dynamics. While transformations to cancel destabilizing terms in the equation could be found, it was not clear how to ensure the stability of the resulting target systems. The crucial step in extending backstepping to these systems involved identifying attainable target systems that incorporated damping terms.
Thus, a problem for a wave equation with anti-damping on the uncontrolled boundary was addressed in [254]. In this scenario, the anti-damping term acted through the boundary condition, making the system unstable with infinitely many eigenvalues in the right-half plane. Another example considers an output-feedback problem with instability at its free end and control on the opposite end [186]. A similar problem of boundary stabilization for a one-dimensional wave equation with an internal spatially varying antidamping term was tackled in [246]. This antidamping term made the uncontrolled system unstable, with all of its eigenvalues located in the right-half plane. The proposed design used a novel backstepping transformation with a 2x2 structure, mapping the unstable system into a stable target system with a desired decay rate. The control gains were expressed in closed form for systems with constant parameters. Other backstepping designs on the wave equation include [315, 295]. Although these backstepping controllers may have limited practical applicability due to their poor delay-robustness margins, resulting from the infinite number of unstable poles in open-loop, the underlying theoretical developments have been essential for the development of the backstepping methodology.
For beams, a pioneering work in this area was [191], where boundary controllers and observers were designed for the wave equation and the Timoshenko beam model with Kelvin-Voigt damping. In contrast to the reaction-diffusion equation, where the potentially destabilizing reaction term can be simply eliminated in some cases, the target systems in [191] contained damping terms arising from Kelvin-Voigt damping (a type of material damping that introduces a strain-rate term in the constitutive equation). This contribution laid the foundation for subsequent studies on the shear beam [185] and the Euler–Bernoulli beam [254].